OpenStudy (anonymous):

Determine the present value, P, you must invest to have the future value, A, at simple interest rate r after time. Round answer to the nearest dollar. A= $196.00 r=10% t=4 years

OpenStudy (anonymous):

P = A / (1+ i ) ^ N P = 196 / ( 1 + .10 ) ^ 4 = 134